OpenStudy (anonymous):

I need to know how to : Find the limit of 3n^3-5n/n^3-2n^2+1 as n approaches infinity. I know the answer is 3 but I don't know how. Please help!! Thank you!

OpenStudy (anonymous):

is it \[\frac{ 3n^3-5n }{n^3-2n^2+1 }\]

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

you can divide... \[\frac{ 3n^3-5n }{ n^3-2n^2+1 }=3+\frac{ 6n^2-3 }{ n^3-2n^2+1 }\] so do you understand that the 2nd term goes to 0 as n goes to infinity?

OpenStudy (anonymous):

that leaves 3.

OpenStudy (anonymous):

What did you divide by? I'm not understanding how you got the numerator.

OpenStudy (anonymous):

divide 3n^3-5n by n^3-2n^2+1... and the numerator should be 6n^2 - 5n - 3

OpenStudy (anonymous):

I got something different but either way it's the same concept so thank you!

OpenStudy (anonymous):

have you learned derivatives yet?

OpenStudy (anonymous):

Nope, we are next chapter.